# SHINE: Social Homology Identification for Navigation in Crowded Environments

### v1

@@ Abstract
Human-Aware Motion Planning, Motion and Path Planning, Integrated Planning and Learning, Collision Avoidance
@@ 2.1 Social navigation
Navigating in dynamic environments poses a challenge that has received significant attention from the robotics community. Early approaches primarily focused on static obstacles, such as the dynamic window approach proposed in Fox et al. (1997), later improved by including moving obstacles in Missura and Bennewitz (2019). Velocity obstacles, introduced in Fiorini and Shiller (1998), laid the foundation for numerous studies, including Wilkie et al. (2009) or Lorente et al. (2018), which used the concept for modeling obstacle motion. Additionally, strategies like the Optimal Reciprocal Collision Avoidance (ORCA) (Van Den Berg et al., 2011) and other methods based on it like NH-ORCA (Alonso-Mora et al., 2012) or VR-ORCA (Guo et al., 2021) have explored the principle of reciprocity in collision avoidance. They consider a shared responsibility between the robot and dynamic obstacles, facilitating collaborative maneuvers. In Lorente et al. (2018), an evolution of the velocity obstacles is presented, where strategies of passing before and after obstacles are applied to reach maximum speed and minimum time to reach the goal.
The motion planners previously mentioned work using a local view of the environment in a short spatial horizon, which may lead to local traps and non-smooth trajectories. In Matínez-Baselga et al. (2023), the planner developed in Lorente et al. (2018) for dynamic environments is integrated with a global planner, avoiding trap situations. Some approaches use a Model Predictive Controller (MPC) to compute optimal control commands for the robot in advance. For example, LMPCC (Brito et al., 2019) is a method for real-time navigation among dynamic obstacles. Crowd motion estimation may be included in the optimizer to make it more reliable (Poddar et al., 2023).
@@
Other approaches use deep learning to teach the robot how to navigate in crowds. Particularly, deep reinforcement learning (DRL) planners show very promising results. A common approach is using decentralized end-to-end algorithms whose output is the velocity of the robot (Chen et al., 2017). Some of the networks use different layers to encode the environment and extract features to navigate, such as LSTM layers (Everett et al., 2018), (Everett et al., 2021), attention (Chen et al., 2019) or graph neural networks (Chen et al., 2020a), (Chen et al., 2020b). However, navigating with humans require special considerations, due to social norms and special needs in collaboration and safety. In Hu et al. (2022), social stress is introduced to the state space of the DRL framework, while in Zhou et al. (2022) and Martinez-Baselga et al. (2023), social attention is used to model the interactions between the humans and the robot. Nevertheless, DRL approaches have the limitations of being very dependent on the simulator and the modeling of people and robot interactions, as well as a lack of explicit collision constraints.
@@ 2.2 Topology-based navigation
Reactive planners focus on the immediate surroundings of the robot, while MPC-based planners optimize their current trajectory, leading to suboptimal paths. Two trajectories with the same start and end points that can not be deformed into the other without intersecting with an obstacle avoid the obstacle by adopting different spatial configurations. We say that they are topological distinct trajectories. Homotopy classes refer to different topological paths or trajectories that are considered equivalent in terms of their spatial relationship (further explained in Section 4). By computing trajectories in distinct homotopy classes, robots can choose trajectories that avoid obstacles and optimize their movements when unexpected obstacles or changes in the environment occur (Kolur et al., 2019). Topological information can be useful not only for navigation, but also for other tasks as information gathering (McCammon and Hollinger, 2021). Discovering topologically distinct trajectories can be done with Voronoi diagrams (Rösmann et al., 2017a), but a more common approach in the literature are Probabilistic Roadmaps (PRM) (Siméon et al., 2000) or derivatives (Novosad et al., 2023), which is a sampling-based method to construct a sparse roadmap.
@@ 3 Problem formulation
where $\bm{x}_{k}\in\mathbb{R}^{n_{x}}$ is the state and $\bm{u}_{k}\in\mathbb{R}^{n_{u}}$ the control input of the system at time step $k$, and $n_{x}$ and $n_{u}$ are the state and input dimensions, respectively. The state of the robot at time $k$ contains its 2-D position $\bm{p}_{k}=(x_{k},y_{k})\in\mathbb{R}^{2}$. We will use the superscript $T$ when referring to the collection of variables over $T$ discrete time steps. For example, $\bm{p}^{T}_{k}=[\bm{p}_{k-T},\dots,\bm{p}_{k}]$ denotes the robot’s positions over the past $T$ discrete time steps.
@@
The state space that we consider for planning is composed by the workspace and time: $\mathcal{X}=\mathbb{R}^{2}\times[0,T]$. A trajectory is a continuous path through the state space: $\bm{\tau}:[0,1]\rightarrow\mathcal{X}$. The collision-free space describes the state space without the space occupied by the obstacles.
@@ 3.1 Homotopy and Homology Classes
We are interested in the high-level behavior of trajectories in the collision-free space (i.e., how they pass the obstacles) so that we can learn this behavior from humans. The direction in which a trajectory passes obstacles is captured by homotopy classes, a concept from topology. Formally, two trajectories are in the same homotopy class if they can be continuously deformed into each other without intersecting with any obstacle, while keeping the endpoints fixed (Bhattacharya et al., 2012). An example is depicted in Fig. 2.
While homotopy classes capture useful information, they are difficult to compute in practice. In this work, we use the almost equivalent concept of homology classes, similar to previous works (Bhattacharya et al., 2012; Rösmann et al., 2017a).
Refer to caption Refer to caption
(a) Top view (b) State space (time in the z-axis)
Figure 2: An example of a two trajectories, left-passing (red line) and right-passing (blue line), in distinct homotopy classes. The trajectories of the robot (blue disk) pass the future motion prediction of the obstacles (transparent cylinders) differently. We introduce a reference trajectory $\bm{\tau}_{ref}$ (dotted yellow) to assess the homotopy class of a single trajectory.
@@
(Bhattacharya et al., 2012) (Homologous Trajectories) Two trajectories, $\bm{\tau}_{1}:[0,1]\rightarrow\mathcal{X}$ and $\bm{\tau}_{2}:[0,1]\rightarrow\mathcal{X}$, connected by the same starting and ending point, are homologous if both $\bm{\tau}_{1}$ and $\bm{\tau}_{2}$ (with $\bm{\tau}_{2}$ in the opposite direction) combine to create the entire boundary of a two-dimensional manifold that is immersed in the complex plane $\mathcal{C}$, and this manifold does not include or intersect any of the obstacles (Bhattacharya et al., 2012). Formally:
$[\bm{\tau}_{1}\sqcup\bm{\tau}_{2}]=\bm{0}\in H_{1}(\mathcal{X}),$ (2)
where $H_{1}(\mathcal{X})$ is the trivial class of the first homology group of $\mathcal{X}$. Homologous trajectories belong to the same homology class.
Contrary to homotopy classes, homology classes are comparable in practice through their H-signature (Bhattacharya et al., 2012). In the 3-dimensional state space, the H-signature can be computed by assuming that the time dimension progresses linearly (e.g., discrete-time). For this computation, a wire (or skeleton) represents obstacles and their future motion through the state space. The obstacle skeleton is looped outside of the state space to create a closed wire loop. Ampere’s Law gives the current passing through the loop $\gamma$:
[…]
where $\bm{B}(\bm{l})$ is the magnetic field vector, $d\bm{l}$ is an infinitesimal vector element along the loop path and $\mu_{0}$ is the permeability of free space (a constant).
For every trajectory $\bm{\tau}$, having $M$ obstacles in the environment and that $C_{1}(\mathbb{R}^{3})$ is the space of all trajectories, its H-signature $\mathcal{H}_{3}:C_{1}(\mathbb{R}^{3})\rightarrow\mathbb{R}^{M}$, is a vector of real numbers:
@@
The H-signature is invariant for every trajectory in a homology class. The homology classes of two trajectories $\bm{\tau}_{1}$ and $\bm{\tau}_{2}$ connected to the same starting and ending points may be compared using their H-signature, by using the loop formed by $\bm{\tau}_{1}\sqcup-\bm{\tau}_{2}$. Using Equation 5 holds that:
$h_{i}(\bm{\tau}_{1}\sqcup-\bm{\tau}_{2})=\int_{\bm{\tau}_{1}\sqcup-\bm{\tau}_{2}}\bm{B}(\bm{l})d\bm{l}$ (6)
Note that the value of $\mu_{0}$ may be set to 1 in Equation 3, and the magnitude of the current in the conductor is assumed to be 1. Therefore, for any closed loop $\gamma$, the value of the integral results 0 iff $\gamma$ does not enclose the conductor and ±$\pm 1$ otherwise. In other words, $h_{i}(\bm{\tau}_{1}\sqcup-\bm{\tau}_{2})$ will be 0 if the trajectories do not enclose the obstacle $i$, and ±$\pm 1$ if they do.
@@ 4.1 Modified H-signature
We define our own version of the H-signature using a reference trajectory $\bm{\tau}_{ref}$ , represented in Figure 2 in yellow, and Equation 6. We use a reference trajectory to be able to assess the topology of a trajectory in isolation, consistently and invariantly. Having a trajectory $\bm{\tau}$ with initial and final times $t_{0}$ and $t_{f}$, whose origin is $\bm{\tau}^{0}=(x_{o},y_{o},t_{0})$ and final point is $\bm{\tau}^{f}=(x_{f},y_{f},t_{N})$; the reference trajectory is defined by the straight line that connects the points $\bm{\tau}^{0}$ and $\bm{\tau}^{f}$.
Among the infinite set of trajectories, that could be chosen, we selected $\bm{\tau}_{ref}$ as the reference path for the signature as it is the trajectory on the minimal energy, and it forms a loop with $\bm{\tau}$. The loop $\bm{\tau}_{ref}\sqcup-\bm{\tau}$ may surround the obstacle $i$, so $h_{i}(\bm{\tau}_{ref}\sqcup-\bm{\tau})$ may have the value 0 if the obstacle is not inside the loop ($\bm{\tau}$ behaves in the same way as $\bm{\tau}_{ref}$), or $\pm$1 otherwise. Since in this paper, we consider $M$ humans as dynamic obstacles, we define the new signature as:
$\mathcal{H}_{ref}(\bm{\tau})=[h_{0}(\bm{\tau}_{ref}\sqcup-\bm{\tau}),\dots,h_{M}(\bm{\tau}_{ref}\sqcup-\bm{\tau})]^{T}$ (7)
In practice, we use the absolute value of the integral as the H-signature, as the sign provides redundant information in this case (the result is 0 if $\bm{\tau}$ does the same as $\bm{\tau}_{ref}$ and 1 otherwise). The benefits of the signature are that it is humanly understandable and invariant of the shape and direction of the trajectories.
@@ 4.2 Homology class selection
We consider a parametric cost function that estimates that difference using as inputs previous observations of the robot’s positions $\bm{p}_{k}^{T}$ and the other humans’ positions in the scenario $\mathcal{O}^{T}_{k}$ over a time horizon $T$ at time $k$. The cost function designed $\mathcal{J}_{\theta}(\mathcal{H}_{ref}(\bm{\tau}^{g}),\bm{p}_{k}^{T},\mathcal{O}^{T}_{k})$ depends on the parameters $\theta$ and is approximated using a deep neural network, represented in a diagram in Figure 3. The goal of the network is to estimate the difference in H-signature (in a Euclidean sense) between the trajectory $\bm{\tau}^{g}$ and the trajectory $\bm{\tau}^{h}$ that a human would choose:

### v3

@@ Abstract
Human-Aware Motion Planning, Motion and Path Planning, Collision Avoidance, Integrated Planning and Learning
@@ 2.1 Social navigation
Navigating in environments populated by humans poses a challenge that has received significant attention from the robotics community, analyzed in recent survey papers as Mavrogiannis et al. (2023) or Singamaneni et al. (2024). Early obstacle avoidance approaches primarily focused on static obstacles, such as the dynamic window approach proposed in Fox et al. (1997), later improved by including moving obstacles in Missura and Bennewitz (2019). Velocity obstacles, introduced in Fiorini and Shiller (1998), laid the foundation for numerous studies, including Wilkie et al. (2009) or Lorente et al. (2018), which used the concept for modeling obstacle motion. Additionally, strategies like the Optimal Reciprocal Collision Avoidance (ORCA) (Van Den Berg et al., 2011) and other methods based on it like NH-ORCA (Alonso-Mora et al., 2012) or VR-ORCA (Guo et al., 2021) have explored the principle of reciprocity in collision avoidance. They consider a shared responsibility between the robot and dynamic obstacles, facilitating collaborative maneuvers. In Lorente et al. (2018), an evolution of the velocity obstacles is presented, where strategies of passing before and after obstacles are applied to reach maximum speed and minimum time to reach the goal.
The motion planners in previously mentioned work use a local view of the environment in a short spatial horizon, which may lead to local traps and non-smooth trajectories. In Matínez-Baselga et al. (2023), the planner developed in Lorente et al. (2018) for dynamic environments is integrated with a global planner, avoiding trap situations. Some approaches use a Model Predictive Controller (MPC) to compute optimal control commands for the robot in advance. For example, LMPCC (Brito et al., 2019) is a method for real-time navigation among dynamic obstacles. Crowd motion estimation may be included in the optimizer to make it more reliable (Poddar et al., 2023).
@@
Previous methods typically treat other agents as dynamic obstacles whose trajectories are predicted, without fully addressing the complexities of interactions between decision-making agents. This may lead to defensive and opaque behaviors (Trautman et al., 2015). Thus, another important body of work addresses multi-agent interactions by coupling prediction and planning. Some methods do it explicitly using, for example, intent inference (Bandyopadhyay et al., 2013; Bai et al., 2015), a game-theoretic perspective (Fridovich-Keil et al., 2020; Sun et al., 2021; Peters et al., 2024; Mehr et al., 2023) or topological representations (Mavrogiannis et al., 2017; Mavrogiannis and Knepper, 2019). Since each agent’s action depends on the decisions of the others, explicitly solving the coupled prediction and planning problem presents a considerable computational challenge.
Other approaches implicitly accounts for interactions with crowds in their design. Particularly, deep reinforcement learning (DRL) planners show promising results by learning from interactions with the environment. A common approach is using decentralized end-to-end algorithms whose output is the velocity of the robot (Chen et al., 2017). Some of the networks use different layers to encode the environment and extract features to navigate, such as LSTM layers (Everett et al., 2018), (Everett et al., 2021), attention (Chen et al., 2019) or graph neural networks (Chen et al., 2020a), (Chen et al., 2020b). However, navigating with humans require special considerations, due to social norms and special needs in collaboration and safety. In Hu et al. (2022), social stress is introduced to the state space of the DRL framework, while in Zhou et al. (2022) and Martinez-Baselga et al. (2023), social attention is used to model the interactions between the humans and the robot. Nevertheless, DRL approaches have the limitations of being very dependent on the simulator and the modeling of people and robot interactions, as well as a lack of explicit collision constraints.
@@ 2.2 Topology-based navigation
Reactive planners focus on the immediate surroundings of the robot, while MPC-based planners optimize their current trajectory, leading to suboptimal paths. Two trajectories that start and end at the same points but cannot be transformed into one another without crossing an obstacle adopt different spatial configurations to avoid the obstacle.We say that they are topological distinct trajectories. Homotopy classes refer to different topological paths or trajectories that are considered equivalent in terms of their spatial relationship (further explained in Section 4). By computing trajectories in distinct homotopy classes, robots can choose trajectories that avoid obstacles and optimize their movements when unexpected obstacles or changes in the environment occur (Kolur et al., 2019). Topological information can be useful not only for navigation, but also for other tasks as information gathering (McCammon and Hollinger, 2021). Discovering topologically distinct trajectories can be done with Voronoi diagrams (Rösmann et al., 2017a), but a more common approach in the literature are Probabilistic Roadmaps (PRM) (Siméon et al., 2000) or derivatives (Novosad et al., 2023), which is a sampling-based method to construct a sparse roadmap.
@@ 3 Problem formulation
where $\bm{x}_{k}\in\mathbb{R}^{n_{x}}$ is the state and $\bm{u}_{k}\in\mathbb{R}^{n_{u}}$ the control input of the system at time step $k$, and $n_{x}$ and $n_{u}$ are the state and input dimensions, respectively. The state of the robot at time $k$ contains its 2-D position $\bm{p}_{k}=(x_{k},y_{k})\in\mathbb{R}^{2}$.
We will use the superscript $T$ when referring to the collection of variables over $T$ discrete time steps. For example, $\bm{p}^{T}_{k}=[\bm{p}_{k-T},\dots,\bm{p}_{k}]$ denotes the robot’s positions over the past $T$ discrete time steps.
@@
The state space that we consider for planning is composed by the workspace and time: $\mathcal{X}=\mathbb{R}^{2}\times[0,T]$. A trajectory is a continuous path through the state space: $\bm{\tau}:[0,1]\rightarrow\mathcal{X}$, and $\mathcal{X}_{traj}$ the set of all possible trajectories in $\mathcal{X}$. We denote $\bm{\tau}_{1}\sqcup\bm{\tau}_{2}$ the the curve formed by $\bm{\tau}_{1}$ and $\bm{\tau}_{2}$ together, and $-\bm{\tau}$ the same trajectory $\bm{\tau}$ but in opposite orientation. The collision-free space describes the state space without the space occupied by the obstacles.
@@ 3.1 Homotopy and Homology Classes
We are interested in the high-level behavior of trajectories in the collision-free space (i.e., how they pass the obstacles) so that we can learn this behavior from humans. The direction in which a trajectory passes obstacles is captured by homotopy classes, a concept from topology.
@@
(Bhattacharya et al., 2012) (Homotopic Trajectories) Two trajectories, $\bm{\tau}_{1}:[0,1]\rightarrow\mathcal{X}$ and $\bm{\tau}_{2}:[0,1]\rightarrow\mathcal{X}$, connected by the same starting and ending point, are homotopic if they can be continuously deformed into each other without intersecting with any obstacle, while keeping the endpoints. Homotopic trajectories belong to the same homotopy class.
(a) Trajectories in different homotopy classes
(b) Trajectories in different homotopy classes and $\bm{\tau}_{ref}$
Figure 2: Schematic representation of trajectories starting at point $\bm{p}_{0}$ and ending at $\bm{p}_{goal}$. Obstacles trajectory is $\mathcal{C}_{i}$. Green and violet trajectories have the same homotopy class among them. 3-D (top) with one axis representing time and 2-D (bottom) representations are plotted in both (a) and (b). In (b) we introduce a reference trajectory $\bm{\tau}_{ref}$ (yellow) to assess the homology class of a single trajectory.
An example of trajectories in two different homotopy classes is depicted in Fig. 2(a). In this work, we use the almost equivalent concept of homology classes, similar to previous works (Bhattacharya et al., 2012; Rösmann et al., 2017a).
Definition 2.
(Bhattacharya et al., 2012) (Homologous Trajectories) Two trajectories, $\bm{\tau}_{1}:[0,1]\rightarrow\mathcal{X}$ and $\bm{\tau}_{2}:[0,1]\rightarrow\mathcal{X}$, connected by the same starting and ending point, are homologous if the loop formed by them ($\bm{\tau}_{1}\sqcup-\bm{\tau}_{2}$) forms the boundary of a two-dimensional manifold that is immersed in $\mathcal{X}$ and does not contain nor intersect with any obstacle. Homologous trajectories belong to the same homology class.
Two homotopic trajectories are always homologous. Even though two homologous trajectories may not be homotopic, they are very similar in most of robotic applications. Particularly, in the state space $\mathcal{X}$, two homologous trajectories are always homotopic (Bhattacharya et al., 2011). In this work, we use the homology classes concept because, contrary to homotopy classes, which are hard to compute, homology classes are comparable in practice through their H-signature (Bhattacharya et al., 2012). In the 3-dimensional state space, the H-signature can be computed by assuming that the time dimension progresses linearly (e.g., discrete-time). To do so, each obstacle $i$ may be defined as a vector field that exhibits rotational symmetry around the time-axis, centered in the 3-dimensional obstacle trajectory $\mathcal{C}_{i}$, as seen in Fig. 2(a). The vector field corresponding to obstacle $i$, $\bm{F}_{i}$, is constructed as an integral over $\mathcal{C}_{i}$:
$\bm{F}_{i}(\bm{r})=\int_{\mathcal{C}_{i}}\frac{d\bm{l}\times(\bm{r}-\bm{l})}{|\bm{r}-\bm{l}|^{3}}$ (2)
where $d\bm{l}$ is an infinitesimal vector element along $\mathcal{C}_{i}$, $\bm{l}$ a point in $\mathcal{C}_{i}$ and $\bm{r}$ the point in which the field is being evaluated. This definition is analogous to the Biot-Savart law of electromagnetism.
An important property of $\bm{F}_{i}$ is that it is not path-independent, i.e., the integral of $\bm{F}_{i}$ between two points depends on the specific path taken, not only on the end-points. Nevertheless, the integral over $\bm{F}_{i}$ of trajectories that belong to the same homology class is the same. In Fig. 2(a), all violet trajectories and all green trajectories exhibit this property, respectively. Thus, path independence holds within homology classes. This is used to define signatures that represent homology classes, named H-signatures. For every trajectory $\bm{\tau}$, having $M$ obstacles in the environment, its H-signature $\mathcal{H}_{3}:\mathcal{X}_{traj}\rightarrow\mathbb{R}^{M}$ is a vector in $\mathbb{R}^{M}$:
@@ 4.1 Modified H-signature
[…]
$h_{i}(\bm{\tau}_{1}\sqcup-\bm{\tau}_{2})=\oint_{\bm{\tau}_{1}\sqcup-\bm{\tau}_{2}}\bm{F}_{i}(\bm{r})d\bm{r}$ (5)
Therefore, $h_{i}(\bm{\tau}_{1}\sqcup-\bm{\tau}_{2})$ will be 0 if the trajectories do not enclose the obstacle $i$, and ±1$\pm$ 1± 1 if they do.
We define our own version of the H-signature using a reference trajectory $\bm{\tau}_{ref}$, represented in Fig. 2(b) in yellow, and Equation 5. We use a reference trajectory to be able to assess the topology of a trajectory in isolation, consistently and invariantly. Having a trajectory $\bm{\tau}$ with initial and final times $t_{0}$ and $t_{f}$, whose origin is $\bm{\tau}^{0}=(x_{o},y_{o},t_{0})$ and final point is $\bm{\tau}^{f}=(x_{f},y_{f},t_{f})$; the reference trajectory is defined by the straight line that connects the points $\bm{\tau}^{0}$ and $\bm{\tau}^{f}$.
Among the infinite set of trajectories, that could be chosen, we selected $\bm{\tau}_{ref}$ as the reference path for the signature as it is the simplest trajectory that drives the robot to the goal if there were not obstacles (a straight line with constant velocity), and it forms a loop with $\bm{\tau}$. We define the new signature for $M$ obstacles as:
$\mathcal{H}_{ref}(\bm{\tau})=[h_{1}(\bm{\tau}_{ref}\sqcup-\bm{\tau}),\dots,h_{M}(\bm{\tau}_{ref}\sqcup-\bm{\tau})]^{T}$ (6)
In practice, and due to our application of $\mathcal{H}_{ref}$, we use the absolute value of the integral as the H-signature, as the sign provides redundant information. In this way, for each obstacle $i$, the element $i$ of $\mathcal{H}_{ref}(\bm{\tau})$ will have the value 0 if $\bm{\tau}$ has the same topological behavior as $\bm{\tau}_{ref}$ with respect to $i$ and $n_{w}$ otherwise (usually 1 in our application, as robot trajectories are normally not supposed to loop around pedestrians). Unlike $\mathcal{H}_{3}$ defined in Equation 3, the meaning of $\mathcal{H}_{ref}$ may be easily understood by humans. Intuitively, $\mathcal{H}_{3}$ encodes how much a trajectory loops around each of the obstacles, with positive or negative values depending on the looping side. Thus, the value of $\mathcal{H}_{3}(\tau)$ depends on the starting and final position of $\tau$ and the specific trajectories of the obstacles, whereas the value of $\mathcal{H}_{ref}$ may be intuitively inferred. For example, $\mathcal{H}_{ref}$ of the violet trajectory in Fig. 2(b) is [1] and $\mathcal{H}_{ref}$ of the green one is [0]. Therefore, the benefits of the signature are that it is humanly understandable and invariant of the shape and direction of the trajectories, as well as to the scenarios and environments.
Every homology class may not have non-colliding and feasible trajectories when considering robot and human radius, specially in crowded environments. For example, Fig. 3 shows a scenario with two humans where the robot may only choose between trajectories in two homology classes. Even though $\bm{\tau}_{ref}$ leads to a collision, $\mathcal{H}_{ref}$ may still be computed.
Refer to caption Refer to caption
(a) Top view (b) State space (time in the z-axis)
Figure 3: An example of two trajectories, left-passing (red line) and right-passing (blue line), in distinct homology classes. The trajectories of the robot (blue disk) pass the future motion prediction of the humans (transparent cylinders) differently. The reference trajectory $\bm{\tau}_{ref}$ (dotted yellow) would lead to collision. Orange dots are other possible goal grid points computed by enlarging the goal with the robot radius.
@@ 4.2 Homology class selection
We consider a parametric cost function that estimates that difference using as inputs previous observations of the robot’s positions $\bm{p}_{k}^{T}$ and the other humans’ positions in the scenario $\mathcal{O}^{T}_{k}$ over a time horizon $T$ at time $k$. The cost function designed $\mathcal{J}_{\theta}(\mathcal{H}_{ref}(\bm{\tau}^{g}),\bm{p}_{k}^{T},\mathcal{O}^{T}_{k})$ is approximated using a deep neural network with parameters $\theta$, represented in a diagram in Fig. 4. The goal of the network is to estimate the difference in H-signature (in a Euclidean sense) between the trajectory $\bm{\tau}^{g}$ and the trajectory $\bm{\tau}^{h}$ that a human would choose: